OpenStudy (anonymous):

help please! which of the folloxing expressions is equivalent to 20-4/5x>16? a. x<5 b. x>5 c. x>32 1/2 d. x< 32 1/2

OpenStudy (anonymous):

All we need to do is simply isolate x. \[20-\frac{ 4 }{ 5 }x>16 \rightarrow -\frac{ 4 }{ 5 }x>-4\] Whenever we multiply or divide by a negative, we must flip the sign of the inequality. \[\frac{ 4 }{ 5 }x<4 \rightarrow \bbox[4px,border:1px solid black]{x<5}\]

OpenStudy (anonymous):

ok thanks I get it :)

OpenStudy (anonymous):

You are welcome :)